# Proper Effective Temperature and Order Parameters in Relativistic Non-Equilibrium Steady States

### v1

@@ I Introduction
However, the external force and the response are not the only parameters that represent the deviation from equilibrium. We may also employ the effective temperature [1] (see e.g. [2] for a review). The effective temperature characterizes the ratio between the fluctuation and the dissipation in NESS, and it agrees with the temperature of the heat bath when the system of interest is in thermal equilibrium. Therefore, the effective temperature (more precisely, the difference between the effective temperature and the heat-bath temperature) also indicates the deviation from equilibrium.
@@
To answer these questions, we analyze the simplest possible setting in the model in the gauge/gravity correspondence [3][4, 5]. Specifically, we study a NESS of a $(2+1)$-dimensional system described by the D3-D5 model [6, 7]. Using the gauge/gravity correspondence, the expectation values of physical quantities, which result from path integrals in the quantum field theory side, can be obtained from classical calculations on the gravity side. In the context of statistical physics, macroscopic physical quantities that should be obtained by coarse-graining the microscopic degrees of freedom are obtained directly from the classical mechanics of gravity. We apply this advantage of the gauge/gravity duality to our analysis of NESSs in the nonlinear region.
The dual field theory of the D3-D5 system is $(3+1)$-dimensional ${\cal N}=4$ large-$N$ $SU(N)$ supersymmetric Yang-Mills (SYM) theory at large ’t Hooft coupling $\lambda$ coupled to a fundamental hypermultiplet on a $(2+1)$-dimensional defect [7]. We consider a single flavor model. We employ the probe approximation that is valid when $N$ is sufficiently greater than the number of the flavors. The field theory is a conformal field theory (CFT) within the probe approximation. When the hypermultiplet is massless, this theory has a $U(1)_{B}\times SU(2)_{V}\times SU(2)_{H}$ global symmetry.
It has been known that the $SU(2)_{V}$ symmetry spontaneously breaks to a $U(1)$ symmetry at low temperatures at a finite density of the global $U(1)_{B}$ charge in the presence of an external magnetic field [8, 9, 10]. At finite temperatures, the phase transition is either first or second order, whereas the system exhibits a Berezinskii–Kosterlitz–Thouless (BKT) transition at zero temperature. To the best of the authors’ knowledge, phase transitions in this system have so far been investigated only for the cases where the D5-brane is static with respect to the D3-brane.
@@
The effective temperature $T_{\rm eff}$ of NESS has also been discussed in the framework of the gauge/gravity duality [11, 12, 13, 14, 15, 16, 17]. $T_{\rm eff}$ is obtained as a Hawking temperature associated with the induced metric on a string worldsheet or the open-string metric on the worldvolume of a D-brane. Further discussion of effective temperatures can be found in [18, 19].
Since our system is relativistic, the effective temperature depends on the frame of the observer. In [20], the proper effective temperature has been introduced and computed. The proper effective temperature is the effective temperature on the rest frame of the particles that participate in the NESS.
@@
Our results also suggest that $T_{\mathrm{eff}}$ includes both the NESS effect and the relativistic effect of motion at velocity $v$. The transformation converting $T_{\mathrm{eff}}$ into $T_{\mathrm{eff}}^{\mathrm{prop}}$ removes this relativistic effect, isolating the pure NESS contribution [20]. In this sense, $T_{\mathrm{eff}}^{\mathrm{prop}}$ is the NESS analog of the relativistic temperature defined by Landsberg [21] and van Kampen [22], while $T_{\mathrm{eff}}$ corresponds to that defined by Einstein [23] and Planck [24].
The organization of the present paper is as follows. In Section II, we review the D3-D5 system in thermal equilibrium. In Section III, we set up the D3-D5 model to realize NESS. Specifically, we consider a system corresponding to a defect moving at a constant velocity in a heat bath. In Section IV, The effective temperature of the NESS is defined holographically. The details of the numerical analysis and the results are presented in Section V. In the last section, we summarize the conclusions and provide related discussions.
II D3-D5 system in thermal equilibrium
@@
$ds^{2}$ $=$ $-u^{-2}\left(1-\frac{u^{4}}{u_{h}^{4}}\right)dt^{2}+u^{-2}(dx^{2}+dy^{2}+dz^{2})$ (1)
$+\frac{1}{u^{2}\left(1-\frac{u^{4}}{u_{h}^{4}}\right)}du^{2}+d\Omega_{5}^{2},$
@@
In the D3-D5 model, we introduce a probe D5-brane into the bulk geometry. We assume that the D5-brane extends along the $(t,x,y,u)$ directions in the AdS-BH, but not in the $z$ direction. The $S^{5}$ part can be divided into two parts. The probe D5-brane wraps an $S^{2}$ part (which we call $S^{2}_{H}$) of the $S^{5}$, whose line element is $\cos\theta\>d\Omega_{2H}$ in (2). The remaining part of the $S^{5}$ is a 3-dimensional subspace, whose metric is given by $d\theta^{2}+\sin^{2}\theta\>d\Omega_{2V}^{2}$. We call the $S^{2}$ part whose line element is given by $\sin\theta\>d\Omega_{2V}$ as $S_{V}^{2}$. Note that the radii of these 2-spheres depend on $\theta$. We employ the static gauge: the worldvolume coordinates of the D5-brane are $(t,x,y,u)$ and the coordinates on $S^{2}_{H}$.
@@
where $T_{D5}$ is the tension of the D5-brane. $\gamma_{ab}$ is the induced worldvolume metric and $F_{ab}=\partial_{a}A_{b}-\partial_{b}A_{a}$ is the field strength of the worldvolume gauge field $A_{a}$. We assume that the worldvolume gauge field takes the form of $A_{a}d\xi^{a}=A_{t}(u)dt+Bxdy$. One may think that $A_{y}$ can have $u$-dependence, but we find that the physical solution for $A_{y}$ does not have $u$-dependence in the present setting. We henceforth use the unit system where $12\pi\alpha^{\prime}=$ 1.
[…]
@@
Here, $g_{\mu\nu}$ is the metric of the background geometry given in (1) and (2). The prime denotes the derivative with respect to $u$. $V_{S^{2}}$ is the volume of unit $S^{2}$ and $V$ is the volume along the $(x,y)$ directions.
@@
where $\sigma$ is the chiral condensate  [^1] We employ the same definition of chiral condensate as in [10], and $\sigma=\frac{1}{2T_{D5}V_{S_{2}}V}\frac{\delta S_{D5}}{\delta m}$. . $m$ is the mass of the hypermultiplet that we set $m=0$ in this work. When $\sigma=0$ the $SU(2)_{V}$ symmetry is preserved whereas it is broken when $\sigma\neq 0$.
@@
III D3-D5 system in NESS
In the previous section, an equilibrium system, i.e. a state in which the defect is static, was outlined. In this section, we consider the case where the defect is in steady motion at a constant velocity $v$ in the $z$ direction. We apply the magnetic field $B$ in the $z$ direction and consider finite charge density $\rho$ on the moving defect. Note that $B$ is invariant under the boost in the $z$ direction. We define $\rho$ as (6) so that it is conjugate to $A_{t}$ on the rest frame of the heat bath: $\rho$ is constant under the boost by definition.
@@ III D3-D5 system in NESS
where $\tilde{z}(u)=0$ at the boundary  [^2] The analysis here is an extension of [30, 31]. See also [32].. Then the induced metric on the D5-brane is written as
@@
$\quad+\left[\frac{1}{u^{2}\left(1-\frac{u^{4}}{u_{h}^{4}}\right)}+u^{-2}{\tilde{z}^{\prime}}{}^{2}+{\theta^{\prime}}^{2}\right]du^{2}$
$\quad+2u^{-2}v\tilde{z}^{\prime}dtdu+\cos^{2}\theta d\Omega_{2H}^{2}.$
@@
${\cal L}_{D5}$ $=-\cos^{2}\!\theta\sqrt{B^{2}\!+\!g_{xx}^{2}}$
$\sqrt{-v^{2}(g_{xx}\theta^{\prime}{}^{2}\!+\!g_{uu}g_{xx})\!-\!g_{tt}(g_{uu}\!+\!g_{xx}\tilde{z}^{\prime}{}^{2}\!+\!{\theta^{\prime}}^{2})\!-\!{A_{t}^{\prime}}^{2}}.$ (11)
@@
where
$K$ $=$ $\left(-g_{tt}-v^{2}g_{xx}\right)\left(B^{2}+g_{xx}^{2}\right)^{2}\left(g_{uu}+\theta^{\prime}{}^{2}\right),$ (16)
@@
We solve the equation of motion for $\theta$ numerically and read the chiral condensate $\sigma$ from the obtained solution. In order to determine the solution, we need to impose appropriate boundary conditions for $\theta(u)$. One of them is $\theta(u)/u=0$ at the boundary $u=0$. The remaining boundary condition is given by (19). To satisfy these two conditions, $F$ must take a specific value: this is how $F$ is determined as a function of the control parameters $(T,B,\rho,v)$.
@@ IV Effective temperature
where $\Gamma=\det(\gamma_{ab}+F_{ab})$. $G_{\mu\nu}$ is the open-string metric [27] given by
@@
When the $SU(2)_{V}$ symmetry is preserved (i.e. when $\theta(u)=0$), we instead consider $\delta w$ as a small fluctuation of the D5-brane. (See Appendix AA for the definition of $\delta w$.) We find that the linearized equation of motion for $\delta w$ is given by
@@
where $g_{ww}$ is given by (33). The fluctuations in (20), (21) and (23) are governed by effective metrics that are given by $(\Gamma/G)^{1/4}\sin\theta\>G_{ab}$ for the symmetry broken phase and by $(\Gamma/G)^{1/4}\sqrt{g_{ww}}G_{ab}$ for the symmetric phase, where $G=\det G_{ab}$.
@@
The effective metric of the D5-brane under consideration exhibits a horizon at the position given by (18), which we refer to as the effective horizon. The Hawking temperature associated with this horizon is computed as
$T_{\rm{eff}}^{2}=\gamma^{-2}\frac{\left(\left(L+{v}^{2}\right)+{\rho}^{2}{u_{*}}^{4}\sec^{4}\theta_{*}+{u_{*}}L{v}^{2}\theta^{\prime}_{*}\tan\theta_{*}\right)}{\pi^{2}{u_{*}}^{2}\left(L+{\rho}^{2}{u_{*}}^{4}\sec^{4}\theta_{*}\right)\left(1+{u_{*}}^{2}{v}^{2}\theta^{\prime}_{*}{}^{2}\right)},$ (24)
where $\gamma=(1-v^{2})^{-1/2}$, $\theta_{*}=\theta(u_{*})$, $\theta^{\prime}_{*}=\theta^{\prime}(u_{*})$, and
$L$ $=$ $1+B^{2}{u_{*}}^{4},$ (25)
$u_{*}$ $=$ $\frac{1}{\pi T}(1-v^{2})^{1/4}.$ (26)
When $\rho=B=0$ and the $SU(2)_{V}$ symmetry is preserved($\theta_{*}=\theta^{\prime}_{*}=0$), we obtain
@@
which agrees with the result of [17].
In [20], it is argued that the proper effective temperature, defined as
@@
characterizes the distribution of the fluctuations in NESSs at finite velocity. The proper effective temperature is the effective temperature measured in the rest frame of the defect in the NESS. By comparing (24) and (28), we observe that the overall $\gamma^{-1}$ factor present in $T_{\rm eff}$ (as given by (24)) is canceled in the definition of $T_{\rm eff}^{\rm prop}$.
V Numerical analysis and result
V.1 Numerical analysis
The equation of motion for $\theta(u)$ contains $A_{t}^{\prime}(u)$ and $\tilde{z}^{\prime}(u)$. The solutions for $A_{t}^{\prime}(u)$ and $\tilde{z}^{\prime}(u)$ are given by (12) and (13), respectively. Then we obtain the equation of motion for $\theta(u)$ in terms of the charge density $\rho$, the magnetic field $B$, and the velocity $v$ by virtue of (19).
Since the equation of motion for $\theta(u)$ is a nonlinear differential equation, we need to solve it numerically. We specify the boundary conditions at $u=u_{*}$. One might think that we need to specify two boundary conditions $\theta(u_{*})$ and $\theta^{\prime}(u_{*})$. However, the contribution of $\theta^{\prime\prime}(u)$ in the equation of motion vanishes in the vicinity of $u=u_{*}$, and the equation of motion becomes first-order differential equation that relates $\theta^{\prime}(u)$ with $\theta(u)$, there. $u=u_{*}$. Therefore, it is enough to specify $\theta(u_{*})$ there.

### v2

# 1-13-27 Kasuga, Bunkyo-ku, Tokyo 112-8551, Japan

@@ 1 Introduction
However, the external force and the response are not the only parameters that represent the deviation from equilibrium. We may also employ the effective temperature Cugliandolo:97 (see e.g. ref. cugliandolo2011effective for a review). The effective temperature characterizes the ratio between the fluctuation and the dissipation in NESS, and it agrees with the temperature of the heat bath when the system of interest is in thermal equilibrium. Therefore, the effective temperature (more precisely, the difference between the effective temperature and the heat-bath temperature) also indicates the deviation from equilibrium.
@@
To answer these questions, we analyze the simplest possible setting in the model in the gauge/gravity correspondence Maldacena:1997re Gubser:1998bc ; Witten:1998qj . Specifically, we study a NESS of a $(2+1)$-dimensional system described by the D3-D5 model Karch:2000gx ; DeWolfe:2001pq . Using the gauge/gravity correspondence, the expectation values of physical quantities, which result from path integrals in the quantum field theory side, can be obtained from classical calculations on the gravity side. In the context of statistical physics, macroscopic physical quantities that should be obtained by coarse-graining the microscopic degrees of freedom are obtained directly from the classical mechanics of gravity. We apply this advantage of the gauge/gravity duality to our analysis of NESSs in the nonlinear region.
The dual field theory of the D3-D5 system is $(3+1)$-dimensional ${\cal N}=4$ large-NN $SU(N)$ supersymmetric Yang-Mills (SYM) theory at large ’t Hooft coupling $\lambda$ coupled to a fundamental hypermultiplet on a $(2+1)$-dimensional defect DeWolfe:2001pq . We consider a single flavor model. We employ the probe approximation that is valid when $N$ is sufficiently greater than the number of the flavors. The field theory is a conformal field theory (CFT) within the probe approximation. When the hypermultiplet is massless, this theory has a $U(1)_{B}\times SU(2)_{V}\times SU(2)_{H}$ global symmetry.
It has been known that the $SU(2)_{V}$ symmetry spontaneously breaks to a $U(1)$ symmetry at low temperatures at a finite density of the global $U(1)_{B}$ charge in the presence of an external magnetic field Jensen:2010ga ; Evans:2010hi ; Evans:2011tk . At finite temperatures, the phase transition is either first or second order, whereas the system exhibits a Berezinskii–Kosterlitz–Thouless (BKT) transition at zero temperature. To the best of the authors’ knowledge, phase transitions in this system have so far been investigated only for the cases where the D5-brane is static with respect to the D3-brane.
@@
The effective temperature $T_{\rm eff}$ of NESS has also been discussed in the framework of the gauge/gravity duality Gubser:2006nz ; Casalderrey-Solana:2007ahi ; Gursoy:2010aa ; Das:2010yw ; Kim:2011qh ; Sonner:2012if ; Nakamura:2013yqa . $T_{\rm eff}$ is obtained as a Hawking temperature associated with the induced metric on a string worldsheet or the open-string metric on the worldvolume of a D-brane. Further discussion of effective temperatures can be found in refs. Hoshino:2014nfa ; Hoshino:2017air .
Since our system is relativistic, the effective temperature depends on the frame of the observer. In ref. Hoshino:2018vne , the proper effective temperature has been introduced and computed. The proper effective temperature is the effective temperature on the rest frame of the particles that participate in the NESS.
@@
Our results also suggest that $T_{\mathrm{eff}}$ includes both the NESS effect and the relativistic effect of motion at velocity $v$. The transformation converting $T_{\mathrm{eff}}$ into $T_{\mathrm{eff}}^{\mathrm{prop}}$ removes this relativistic effect, isolating the pure NESS contribution Hoshino:2018vne . In this sense, $T_{\mathrm{eff}}^{\mathrm{prop}}$ is the NESS analog of the relativistic temperature defined by Landsberg Landsberg_1966 and van Kampen vanKampen1968 , while $T_{\mathrm{eff}}$ corresponds to that defined by Einstein einstein1907uber and Planck planck1907 .
The organization of the present paper is as follows. In section 2, we review the D3-D5 system in thermal equilibrium. In section 3, we set up the D3-D5 model to realize NESS. Specifically, we consider a system corresponding to a defect moving at a constant velocity in a heat bath. In section 4, the effective temperature of the NESS is defined holographically. The details of the numerical analysis and the results are presented in section 5. In the last section, we summarize the conclusions and provide related discussions.
2 D3-D5 system in thermal equilibrium
@@
$ds^{2}=-u^{-2}\left(1-\frac{u^{4}}{u_{h}^{4}}\right)dt^{2}+u^{-2}(dx^{2}+dy^{2}+dz^{2})+\frac{1}{u^{2}\left(1-\frac{u^{4}}{u_{h}^{4}}\right)}du^{2}+d\Omega_{5}^{2},$ (1)
@@
In the D3-D5 model, we introduce a probe D5-brane into the bulk geometry. We assume that the D5-brane extends along the $(t,x,y,u)$ directions in the AdS-BH, but not in the $z$ direction. The $S^{5}$ part can be divided into two parts. The probe D5-brane wraps an $S^{2}$ part (which we call $S^{2}_{H}$) of the $S^{5}$, whose line element is $\cos\theta\>d\Omega_{2H}$ in eq. (2). The remaining part of the $S^{5}$ is a 3-dimensional subspace, whose metric is given by $d\theta^{2}+\sin^{2}\theta\>d\Omega_{2V}^{2}$. We call the $S^{2}$ part whose line element is given by $\sin\theta\>d\Omega_{2V}$ as $S_{V}^{2}$. Note that the radii of these 2-spheres depend on $\theta$. We employ the static gauge: the worldvolume coordinates of the D5-brane are $(t,x,y,u)$ and the coordinates on $S^{2}_{H}$.
@@
where $T_{D5}$ is the tension of the D5-brane. $\gamma_{ab}$ is the induced worldvolume metric and $F_{ab}=\partial_{a}A_{b}-\partial_{b}A_{a}$ is the field strength of the worldvolume gauge field $A_{a}$. We assume that the worldvolume gauge field takes the form of $A_{a}d\xi^{a}=A_{t}(u)dt+Bxdy$. One may think that $A_{y}$ can have $u$-dependence, but we find that the physical solution for $A_{y}$ does not have $u$-dependence in the present setting. We henceforth use the unit system where $2\pi\alpha^{\prime}=1$.
Now, the D5-brane action is given by $S_{D5}=T_{D5}V_{S^{2}}V\int dtdu{\cal L}_{D5}$, where
@@
Here, $g_{\mu\nu}$ is the metric of the background geometry given in eqs. (1) and (2). The prime denotes the derivative with respect to $u$. $V_{S^{2}}$ is the volume of unit $S^{2}$ and $V$ is the volume along the $(x,y)$ directions.
@@
where $\sigma$ is the chiral condensate. [^1] We employ the same definition of chiral condensate as in ref. Evans:2011tk , and $\sigma=\frac{1}{2T_{D5}V_{S_{2}}V}\frac{\delta S_{D5}}{\delta m}$. $m$ is the mass of the hypermultiplet that we set $m=0$ in this work. When $\sigma=0$ the $SU(2)_{V}$ symmetry is preserved whereas it is broken when $\sigma\neq 0$.
@@
3 D3-D5 system in NESS
In the previous section, an equilibrium system, i.e. a state in which the defect is static, was outlined. In this section, we consider the case where the defect is in steady motion at a constant velocity $v$ in the $z$ direction. We apply the magnetic field $B$ in the $z$ direction and consider finite charge density $\rho$ on the moving defect. Note that $B$ is invariant under the boost in the $z$ direction. We define $\rho$ as eq. (6) so that it is conjugate to $A_{t}$ on the rest frame of the heat bath: $\rho$ is constant under the boost by definition.
@@ 3 D3-D5 system in NESS
where $\tilde{z}(u)=0$ at the boundary. [^2] The analysis here is an extension of refs. Herzog:2006gh ; Gubser:2006bz . See also ref. Karch:2007pd . Then the induced metric on the D5-brane is written as
@@
$\quad+\left[\frac{1}{u^{2}\left(1-\frac{u^{4}}{u_{h}^{4}}\right)}+u^{-2}{\tilde{z}^{\prime}}{}^{2}+{\theta^{\prime}}^{2}\right]du^{2}+2u^{-2}v\tilde{z}^{\prime}dtdu+\cos^{2}\theta d\Omega_{2H}^{2}.$
@@
${\cal L}_{D5}=-\cos^{2}\!\theta\sqrt{B^{2}\!+\!g_{xx}^{2}}\sqrt{-v^{2}(g_{xx}\theta^{\prime}{}^{2}\!+\!g_{uu}g_{xx})\!-\!g_{tt}(g_{uu}\!+\!g_{xx}\tilde{z}^{\prime}{}^{2}\!+\!{\theta^{\prime}}^{2})\!-\!{A_{t}^{\prime}}^{2}}.$ (11)
@@
We solve the equation of motion for $\theta$ numerically and read the chiral condensate $\sigma$ from the obtained solution. In order to determine the solution, we need to impose appropriate boundary conditions for $\theta(u)$. One of them is $\theta(u)/u=0$ at the boundary $u=0$. The remaining boundary condition is given by eq. (19). To satisfy these two conditions, $F$ must take a specific value: this is how $F$ is determined as a function of the control parameters $(T,B,\rho,v)$.
@@ 4 Effective temperature
where $\Gamma=\det(\gamma_{ab}+F_{ab})$. $G_{\mu\nu}$ is the open-string metric Seiberg:1999vs given by
@@
When the $SU(2)_{V}$ symmetry is preserved (i.e. when $\theta(u)=0$), we instead consider $\delta w$ as a small fluctuation of the D5-brane. (See appendix A for the definition of $\delta w$.) We find that the linearized equation of motion for $\delta w$ is given by
@@
where $g_{ww}$ is given by eq. (31). The fluctuations in eqs. (20), (21) and (23) are governed by effective metrics that are given by $(\Gamma/G)^{1/4}\sin\theta\>G_{ab}$ for the symmetry broken phase and by $(\Gamma/G)^{1/4}\sqrt{g_{ww}}G_{ab}$ for the symmetric phase, where $G=\det G_{ab}$.
@@
The effective metric of the D5-brane under consideration exhibits a horizon at the position given by eq. (18), which we refer to as the effective horizon. The Hawking temperature associated with this horizon is computed as
$T_{\rm{eff}}^{2}=\gamma^{-2}\frac{\left(\left(1+B^{2}{u_{*}}^{4}+{v}^{2}\right)+{\rho}^{2}{u_{*}}^{4}\sec^{4}\theta_{*}+{u_{*}}(1+B^{2}{u_{*}}^{4}){v}^{2}\theta^{\prime}_{*}\tan\theta_{*}\right)}{\pi^{2}{u_{*}}^{2}\left(1+B^{2}{u_{*}}^{4}+{\rho}^{2}{u_{*}}^{4}\sec^{4}\theta_{*}\right)\left(1+{u_{*}}^{2}{v}^{2}\theta^{\prime}_{*}{}^{2}\right)},$ (24)
where $\gamma=(1-v^{2})^{-1/2}$, $\theta_{*}=\theta(u_{*})$, $\theta^{\prime}_{*}=\theta^{\prime}(u_{*})$, and $u_{*}=(1-v^{2})^{1/4}/\pi T$. When $\rho=B=0$ and the $SU(2)_{V}$ symmetry is preserved($\theta_{*}=\theta^{\prime}_{*}=0$), we obtain
@@
which agrees with the result of ref. Nakamura:2013yqa .
In ref. Hoshino:2018vne , it is argued that the proper effective temperature, defined as
@@
characterizes the distribution of the fluctuations in NESSs at finite velocity. The proper effective temperature is the effective temperature measured in the rest frame of the defect in the NESS. By comparing eqs. (24) and (26), we observe that the overall $\gamma^{-1}$ factor present in $T_{\rm eff}$ (as given by eq. (24)) is canceled in the definition of $T_{\rm eff}^{\rm prop}$.
5 Numerical analysis and result
5.1 Numerical analysis
The equation of motion for $\theta(u)$ contains $A_{t}^{\prime}(u)$ and $\tilde{z}^{\prime}(u)$. The solutions for $A_{t}^{\prime}(u)$ and $\tilde{z}^{\prime}(u)$ are given by eqs. (12) and (13), respectively. Then we obtain the equation of motion for $\theta(u)$ in terms of the charge density $\rho$, the magnetic field $B$, and the velocity $v$ by virtue of eq. (19).
Since the equation of motion for $\theta(u)$ is a nonlinear differential equation, we need to solve it numerically. We specify the boundary conditions at $u=u_{*}$. One might think that we need to specify two boundary conditions $\theta(u_{*})$ and $\theta^{\prime}(u_{*})$. However, the contribution of $\theta^{\prime\prime}(u)$ in the equation of motion vanishes in the vicinity of $u=u_{*}$, and the equation of motion becomes first-order differential equation that relates $\theta^{\prime}(u)$ with $\theta(u)$, there. Therefore, it is enough to specify $\theta(u_{*})$ there.